Three point charges, q1 = 2.0 μC, q2 = 2.0 μC, and q3 = −1.0 μC, are located at the vertices of an equilateral triangle of side length 30 cm. What is the potential energy of this arrangement?
1
Expert's answer
2020-11-06T10:12:45-0500

"W=\frac kr(q_1q_2+q_1q_3+q_2q_3)=\frac{9\cdot10^9}{0.3}\cdot (2\cdot2-2\cdot1-2\cdot1)\cdot10^{-12}=0."